Find out the value of cot6°× cot66°× cot42°× cot78° such as the given value is equal to 1.

Grade:11

1 Answers

Vikas TU
7 years ago
Here use the identities,
sinA*sin(A+60)*sin(60-A) = sin(3A)/4 and cosA*cos(A+60)*cos(60-A) = cos(3A)/4
Now divide the costerms eqn with sin terms to form the cot relation.
therefore, 
cotA*cot(A+60)*cot(60-A) = cot(3A)
Starting with A = 60 degree taking and by adjusting the cot and tan terms by multiplying and dividing U would get 1 as the final answer.